# TSkips: Efficiency Through Explicit Temporal Delay Connections in Spiking Neural Networks

### v1

@@ 3.4 Exploring the search space
Identifying optimal models with TSkips is challenging due to the vast number of potential architectures and associated TSkip configurations. To manage this complexity, we initialize our search with a backbone network [5]. This backbone provides a starting point for exploration, reducing the number of possible configurations and allowing the search to focus on optimizing the specific parameters of our proposed method. However, significant complexity remains due to the numerous choices for $\Delta t$, connection placement, and network depth.
To illustrate this complexity, consider a 3-layer network with the possibility of forward and backward connections between any two layers (Fig. 4). This simple example results in $12$ possible combinations of skip connections. With a sequence length of $T=6$, this grows to $120$ possible delay combinations and $2^{120}$ total possible TSkip configurations. This combinatorial problem, grows exponentially with the sequence lengths, as each additional time step introduces new TSkip configurations.
Figure 4: Illustration of the NAS Search Space. (a) Example of possible TSkips combinations in a 3-layer network with $\Delta t=6$. (b) Top-ranked forward and backward TSkips with $\Delta t=4$ and $\Delta t=3$, respectively, found by [27]. For visual clarity, the figure depicts the TSkips spanning a single delay window $(\Delta t)$.
This large search space necessitates an efficient exploration method to identify optimal architectures. To address this, we leverage NASWOT-SAHD [27], a training-free NAS method tailored for SNNs. NASWOT-SAHD scores networks at initialization using the Sparsity Aware Hamming Distance (SAHD), which quantifies the dissimilarity in spike patterns between different layers. By favoring networks with diverse spiking activity, SAHD acts as a proxy for identifying optimal sub networks of the backbone architecture augmented with TSkips. Specifically, [27] performs a random search for architectures with optimal $\Delta t$, TSkip positions (origin and destination) and network depth.
@@ 3.5 Training with Explicit Temporal Delays
Training SNNs and hybrid models with gradient descent requires addressing the non-differentiable nature of spike activations [11] in LIF neurons. We employ adaptive LIF neurons [29] with learnable leaks and thresholds, and use the ArcTangent [11] surrogate gradient to approximate the derivative of the spike activation function. To capture precise spike timing and spatio-temporal dependencies, we unroll the network in time. This, combined with adaptive LIF neurons and TSkips, allows for explicit representation of the network’s temporal evolution, capturing both long-term and local temporal patterns. We then use back propagation through time (BPTT) [53] for gradient propagation and Batch Normalization Through Time (BNTT) [26] for improved training stability.
@@ 4.1 Experimental Setup
The DSEC-Flow dataset [13] consists of 24 challenging driving sequences with varying lighting conditions, fast motion, and occlusions. To assess our models on unseen data, we created a custom test set by splitting the training set, following the approach in [41]. To evaluate the accuracy of the predicted optical flow, we use average end point error (AEE) [58], which measures the average Euclidean distance between the predicted and ground truth flow vectors.
@@
For experiments on DSEC-flow [13], we use a fully spiking [29] and hybrid [36] multi-scale encoder-decoder network inspired by EV-FlowNet [58] as our backbone architecture. We augment this architecture by integrating TSkips into the existing skip connections between the encoder and decoder blocks, as depicted in Fig. 5. Due to the inherent structure of our backbone, the search space for optimal TSkips configurations is relatively small, focusing primarily on selecting which existing skip connection to modify and determining the appropriate temporal delay ($\Delta t$).
Figure 5: Fully spiking EV-FlowNet [58] architecture with forward and backward TSkips. Each TSkip replaces a pre-existing skip connection between the encoder and decoder. Although shown together for conciseness, forward and backward connections are evaluated independently. In practice, when used together, they connect different encoder-decoder layers. This architecture can be modified to be a hybrid ANN-SNN model, by replacing the LIF neurons with ReLU activations.
@@
The IBM DVS128 Gesture dataset [21] contains 1342 instances of 11 hand gestures, captured with a DVS128 [33] camera. For experiments on DVS128 Gesture, we use a ResNet18 [17] backbone architecture [5] and explore various configurations by modifying its basic blocks. This search includes modifying the input and output channels of each layer, kernel sizes, stride, and network depth. To integrate TSkips in the backbone, we remove the original ResNet18 skip connections and search for optimal TSkip placement and corresponding $\Delta t$.
For optical flow estimation and gesture recognition, we introduce a learnable scaling factor, $\alpha$, within the TSkip to further enhance temporal processing (detailed in Section. 3.3).
@@
The SHD [8] dataset comprises 10,000 recordings of spoken digits (zero to nine in English and German), while the larger SSC [8] dataset contains 100,000 recordings of spoken words from various speakers. These datasets present challenges for audio classification, requiring the network to capture and process subtle temporal patterns within spike trains that encode spoken sounds. We focus on the top-one classification task for all 35 classes in SSC and all 20 classes in SHD. For our experiments on speech recognition, we use a fully connected Multi Layer Perceptron(MLP) [47] as our backbone architecture. We explore various MLP configurations by adjusting the input and output features of each layer, the overall network depth, and the placement and $\Delta t$ of TSkips. In these experiments, we omit $\alpha$ used in the EV-FlowNet and DVS128 Gesture TSkip architectures.
To improve generalization, we augment the training sets for DVS128 Gesture, SHD and SSC with channel jitter and random noise [49]. We use sequence lengths (T) of 10, 30, and 99 for DSEC-flow, DVS128 Gesture, and SHD/SSC, respectively. All models are trained on an NVIDIA A40 GPU using the Adam optimizer [28]. For DSEC-flow, we use a multi-step learning rate scheduler (initial rate: $10^{-3}$, scaled by 0.7 every 10 epochs) and the supervised mean squared error (MSE) loss [29, 36] for 200 epochs. For all other datasets, we use a Cosine Annealing Learning Rate Scheduler [34] (initial rate: 0.001, minimum rate: $5\times 10^{-6}$, updated every 10 iterations) for 100 epochs.
Details on the constraints on the search parameters and TSkip configurations for DSEC-flow, DVS128 Gesture, SHD, and SSC can be found in the Supplementary material.
@@ 4.2 Ablation Study
Figure 6: Ablation study results on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkip position and (c) network depth on classification accuracy.
As detailed in Section 3.4, the search space of possible TSkips configurations is exponentially large. To effectively navigate this complex search space, we performed an ablation study on an 8-layer MLP baseline found using [27] on the SHD dataset. We varied the temporal delay ($\Delta t$), TSkip position and network depth using concatenation-based [22] TSkips for our analysis (see Supplementary for detailed baseline network configuration and comparisons with addition-based TSkips).
@@
Impact of TSkip Position:
Next, we investigated the impact of TSkip position on classification accuracy (Fig. 6(b)). Fixing $\Delta t=16$, we varied the destination layer of both forward and backward connections. Forward connections originated from the input data, while backward connections originated from the final output. Our analysis revealed that connecting TSkips to deeper layers consistently improved accuracy. This suggests that propagating temporal information to deeper layers is crucial for enhancing model performance. Note, the optimal placement varies based on forward or backward TSkip.
@@
These findings underscore the complex interplay between $\Delta t$, TSkip position (origin and destination layers), and network depth, which we set as our search parameters. For the results presented in the following sections, we leveraged [27] to identify optimal baseline and concatenation-based [22] TSkip architectures, constraining the architecture search by the number of parameters.
@@ 4.3 DSEC-Flow Results
Incorporating TSkips into the fully spiking and hybrid EV-FlowNet architectures [29, 36], as detailed in Section 4.1, significantly lowers AEE on the DSEC-flow dataset (Table 1). Across both fully spiking and hybrid architectures, TSkips consistently yielded an average reduction in AEE of 14% and 9.5%, respectively. Notably, these improvements are achieved with models 3×3$\times$3 × smaller than baselines with comparable AEE. While SOTA methods like E-RAFT [14] and E-FlowFormer [32] achieve 6% and 10% lower AEE, respectively, these reductions come at the cost of increased complexity. E-RAFT, for example, while being 3×3$\times$3 × smaller than our best model, relies on complex mechanisms such as gated recurrent units and iterative flow refinement. Similalry, E-FlowFormer employs a computationally expensive transformer architecture and is trained on both a custom dataset and DSEC-flow. These complexities likely lead to higher inference energy compared to our lightweight TSkip architectures. Importantly, our approach achieves comparable qualitative performance to baselines 3×3$\times$3 × larger (Fig. 7), highlighting its efficiency in enabling accurate temporal processing without increasing model complexity. Hyper parameters and a detailed analysis of inference energy for TSkips on DSEC-flow are in the Supplementary material.
(a) GT
(b) Mini - Baseline
(c) Mini - F Tskip
(d) Mini - B Tskip
(e) Base - Baseline
Figure 7: Qualitative results on the DSEC-flow dataset. (a) Ground Truth (GT) mask, (b) Mini (3.4M) SNN baseline model mask, (c) Mini SNN (3.4M) with forward (F) Tskip, (d) Mini SNN (3.4M) with backward (B) TSkip, and (e) Base (13M) SNN baseline model mask. As observed, the Mini model that is 3.8$\times$ smaller than Base model performs just as well qualitatively when TSkips are incorporated.
Table 1: AEE on DSEC-flow (lower is better). Best performing SNN and hybrid baselines and forward(F)/backward(B) TSkip models are highlighted in bold.
@@
EV-FlowNet [14] 13.04 2.32 -
LSTM-FlowNet [41] N/A [^1] N/A - not reported #Params 1.28 -
E-RAFT [14] 5.3 0.79 -
E-FlowFormer [32] N/A 0.76 -
@@
Base + F TSkip 1.12 0.86
Base + B TSkip 1.13 0.84
@@
Mini + F TSkip 1.44 1.02
Mini + B TSkip 1.46 1.01
@@
Micro + F TSkip 1.57 1.12
Micro + B TSkip 1.56 1.09
@@
Nano + F TSkip 1.86 1.37
Nano + B TSkip 1.77 1.35
@@
Pico + F TSkip 2.19 1.78
Pico + B TSkip 2.28 1.81
Table 2: Classification accuracy on DVS-Gesture. Comparisons between forward(F)/backward(B) TSkips and SOTA models. Best models in bold.
@@
SLAYER [50] SNN (8 layer) N/A 93.6
KLIF [23] VGG-11 N/A 93.75
DECOLE [24] Custom N/A 95.2
SNN-Skip [2] ResNet18 N/A 95.43
SCRNN [54] Custom N/A 96.59
STS-ResNet [48] ResNet18 N/A 96.7
Mamba-Spike [42] Mambda 6.1 97.8
@@
vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. 0.30 89.12
F TSkip 0.42 94.82
B TSkip 0.38 95.97
F + B TSkip 0.55 97.52
4.4 DVS128-Gesture, SHD and SSC Results
On the DVS Gesture dataset, incorporating TSkips resulted in a significant 8.37% improvement in accuracy over the baseline model, a convolutional SNN found by [27]. Importantly, our approach achieved this with significantly smaller models compared to SOTA, including a spiking transformer [42], demonstrating the efficiency of our proposed method. Furthermore, our method outperformed several other approaches, including spiking RNNs [54], methods incorporating learnable delays [50], and methods using specialized LIF neurons or spatio-temporal feature extraction [23, 48]. This highlights the effectiveness of our approach in capturing and leveraging temporal dependencies for improved gesture recognition.
Table 3: Classification accuracy on SHD and SSC datasets. Comparisons between forward(F), backward(B) and a combination (F+B) of both TSkips against SOTA models and two baseline models: a shallow 4-layer MLP network (Baseline - 1) and a deep 8-layer MLP (Baseline - 2). Best models with a single and two TSkips in bold.
@@ 4.4 DVS128-Gesture, SHD and SSC Results
LSTM-LIF [56] 0.14 0.11 88.91 63.46
SRNN [55] N/A N/A 90.4 74.2
DL256-SNN-DLoss [51] 0.14 - 92.56 -
SpikGRU [9] - 0.28 - 77
radLIF [3] 3.9 3.9 94.62 77.40
Spikingformer [57]333[57] was evaluated on SHD and SSC. [57] for learning rate, with minimal updates to the architecture. 1.98 1.99 82.68 72.43
DCLS-Delays [52] 0.21 2.5 95.07 80.69
@@
vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. - 1 0.16 0.12 68.5 71.2
F TSkip - 1 0.24 0.42 92.32 76.5
B TSkip - 1 0.19 0.24 93.64 79.87
F + B TSkip - 1 0.20 0.54 93.01 78.64
vRNN [^2] Backward skips are added to all baseline layers with $\Delta~{}t=1$. - 2 1.04 1.04 71.35 72.24
F TSkip - 2 1.12 1.08 94.15 78.98
B TSkip - 2 1.16 1.14 93.86 79.65
F + B TSkip - 2 1.28 1.42 94.73 80.23
@@
While TSkips achieves accuracy within 0.34% of SOTA model DCLS-Delays [16] on SHD and 0.46% on SSC, it offers several advantages: applicability across various network architectures, as shown in Section 4.3, efficiency due to very few new trainable parameters, and strong performance. Notably TSkips achieves comparable accuracy to [16] on the SSC dataset with a 44% smaller model.

### v2

@@ 3.4 Exploring the search space
Figure 4: Illustration of the NAS Search Space. (a) Example of possible TSkips combinations in a 3-layer network with $T=6$. (b) Top-ranked forward and backward TSkips with $\Delta t=4$ and $\Delta t=3$, respectively, found by (Kim et al., 2022). For visual clarity, the figure depicts the TSkips spanning a single delay window $(\Delta t)$. Here $t_{0}$ refers to the first time step and $T$ is the sequence length.
Identifying optimal models with TSkips is challenging due to the vast number of potential architectures and associated TSkips configurations. To manage this complexity, we initialize our search with a backbone network (Cai et al., 2020). This backbone provides a starting point for exploration, reducing the number of possible configurations and allowing the search to focus on optimizing the specific parameters of our proposed method. However, significant complexity remains due to the numerous choices for $\Delta t$, connection placement, and network depth.
To illustrate this complexity, consider a 3-layer network with the possibility of forward and backward connections between any two layers (Fig. 4). This simple example results in $12$ possible combinations of skip connections. With a sequence length of $T=6$, this grows to $120$ possible delay combinations and $2^{120}$ total possible TSkips configurations. This combinatorial problem, grows exponentially with the sequence lengths, as each additional time step introduces new TSkips configurations.
This large search space necessitates an efficient exploration method to identify optimal architectures. To address this, we leverage NASWOT-SAHD (Kim et al., 2022), a training-free NAS method tailored for SNNs. NASWOT-SAHD scores networks at initialization using the Sparsity Aware Hamming Distance (SAHD), which quantifies the dissimilarity in spike patterns between different layers. By favoring networks with diverse spiking activity, SAHD acts as a proxy for identifying optimal sub networks of the backbone architecture augmented with TSkips. Specifically, (Kim et al., 2022) performs a random search for architectures with optimal $\Delta t$, TSkips positions (origin and destination) and network depth.
@@ 3.5 Training with Explicit Temporal Delays
Training SNNs and hybrid models with gradient descent requires addressing the non-differentiable nature of spike activations (Fang et al., 2021a) in LIF neurons. We employ adaptive LIF neurons (Kosta & Roy, 2023) with learnable leaks and thresholds, and use the ArcTangent (Fang et al., 2021a) surrogate gradient to approximate the derivative of the spike activation function. To capture precise spike timing and spatio-temporal dependencies, we unroll the network in time. This, combined with adaptive LIF neurons and TSkips, allows for explicit representation of the network’s temporal evolution, capturing both long-term and local temporal patterns. We then use back propagation through time (BPTT) (Werbos, 1990) for gradient propagation and Batch Normalization Through Time (BNTT) (Kim & Panda, 2021) for improved training stability.
@@ 4.1 Experimental Setup
The DSEC-Flow dataset (Gehrig et al., 2021a) consists of 24 challenging driving sequences with varying lighting conditions, fast motion, and occlusions. To assess our models on unseen data, we created a custom test set by splitting the training set, following the approach in Ponghiran et al. (2023). To evaluate the accuracy of the predicted optical flow, we use average end point error (AEE) (Zhu et al., 2018a), which measures the average Euclidean distance between the predicted and ground truth flow vectors.
@@
For experiments on DSEC-flow (Gehrig et al., 2021a), we use a fully spiking (Kosta & Roy, 2023) and hybrid (Negi et al., 2024) multi-scale encoder-decoder network inspired by EV-FlowNet (Zhu et al., 2018a) as our backbone architecture. We augment this architecture by integrating TSkips into the existing skip connections between the encoder and decoder blocks, as depicted in Fig. 5. Due to the inherent structure of our backbone, the search space for optimal TSkips configurations is relatively small, focusing primarily on selecting which existing skip connection to modify and determining the appropriate temporal delay ($\Delta t$).
Figure 5: Fully spiking EV-FlowNet (Zhu et al., 2018a) architecture with forward and backward TSkips. Each TSkips replaces a pre-existing skip connection between the encoder and decoder. Although shown together for conciseness, forward and backward connections are evaluated independently. In practice, when used together, they connect different encoder-decoder layers. This architecture can be modified to be a hybrid ANN-SNN model, by replacing the LIF neurons with ReLU activations.
@@
The IBM DVS128 Gesture dataset (Hu et al., 2022) contains 1342 instances of 11 hand gestures, captured with a DVS128 (Lichtsteiner et al., 2008) camera. For experiments on DVS128 Gesture, we use a ResNet18 (He et al., 2015) backbone architecture (Cai et al., 2020) and explore various configurations by modifying its basic blocks. This search includes modifying the input and output channels of each layer, kernel sizes, stride, and network depth. To integrate TSkips in the backbone, we remove the original ResNet18 skip connections and search for optimal TSkips placement and corresponding $\Delta t$.
For optical flow estimation and gesture recognition, we introduce a learnable scaling factor, $\alpha$, within the TSkips to further enhance temporal processing (detailed in Section. 3.3).
@@
The SHD (Cramer et al., 2022) dataset comprises 10,000 recordings of spoken digits (zero to nine in English and German), while the larger SSC (Cramer et al., 2022) dataset contains 100,000 recordings of spoken words from various speakers. These datasets present challenges for audio classification, requiring the network to capture and process subtle temporal patterns within spike trains that encode spoken sounds. We focus on the top-one classification task for all 35 classes in SSC and all 20 classes in SHD. For our experiments on speech recognition, we use a fully connected Multi Layer Perceptron(MLP) (Rumelhart et al., 1986) as our backbone architecture. We explore various MLP configurations by adjusting the input and output features of each layer, the overall network depth, and the placement and $\Delta t$ of TSkips. In these experiments, we omit $\alpha$ used in the EV-FlowNet and DVS128 Gesture TSkips architectures.
To improve generalization, we augment the training sets for DVS128 Gesture, SHD and SSC with channel jitter and random noise (Shen et al., 2023). We use sequence lengths (T) of 10, 30, and 99 for DSEC-flow, DVS128 Gesture, and SHD/SSC, respectively. All models are trained on an NVIDIA A40 GPU using the Adam optimizer (Kingma & Ba, 2014). For DSEC-flow, we use a multi-step learning rate scheduler (initial rate: $10^{-3}$, scaled by 0.7 every 10 epochs) and the supervised mean squared error (MSE) loss (Kosta & Roy, 2023; Negi et al., 2024) for 200 epochs. For all other datasets, we use a Cosine Annealing Learning Rate Scheduler (Loshchilov & Hutter, 2017) (initial rate: 0.001, minimum rate: $5\times 10^{-6}$, updated every 10 iterations) for 100 epochs.
Details on the constraints on the search parameters and TSkips configurations for DSEC-flow, DVS128 Gesture, SHD, and SSC can be found in Appendix D.1.
@@ 4.2 Ablation Study
Figure 6: Ablation study results on the SHD dataset demonstrating the impact of varying the (a) temporal delay $(\Delta t)$, (b) TSkips position and (c) network depth on classification accuracy.
As detailed in Section 3.4, the search space of possible TSkips configurations is exponentially large. To effectively navigate this complex search space, we performed an ablation study on an 8-layer MLP baseline found using Kim et al. (2022) on the SHD dataset. We varied the temporal delay ($\Delta t$), TSkips position and network depth using concatenation-based (Huang et al., 2017) TSkips for our analysis (see Appendix C.6 for detailed baseline network configuration).
@@
Impact of TSkips Position:
Next, we investigated the impact of TSkips position on classification accuracy (Fig. 6(b)). Fixing $\Delta t=16$, we varied the destination layer of both forward and backward connections. Forward connections originated from the input data, while backward connections originated from the final output. Our analysis revealed that connecting TSkips to deeper layers consistently improved accuracy. This suggests that propagating temporal information to deeper layers is crucial for enhancing model performance. Note, the optimal placement varies based on forward or backward TSkips.
@@
These findings underscore the complex interplay between $\Delta t$, TSkips position (origin and destination layers), and network depth, which we set as our search parameters. For the results presented in the following sections, we leveraged Kim et al. (2022) to identify optimal baseline and concatenation-based (Huang et al., 2017) TSkips architectures (see Appendix C.5 for comparisons with addition-based TSkips), constraining the architecture search by the number of parameters.
@@ 4.3 DSEC-Flow Results
Figure 7: Qualitative results on the DSEC-flow dataset. (a) Ground Truth (GT) mask, (b) Mini (3.4M) SNN baseline model mask, (c) Mini SNN (3.4M) with forward (F) TSkips, (d) Mini SNN (3.4M) with backward (B) TSkips, and (e) Base (13M) SNN baseline model mask. As observed, the Mini model that is 3.8$\times$ smaller than Base model performs just as well qualitatively when TSkips are incorporated.
Incorporating TSkips into the fully spiking and hybrid EV-FlowNet architectures (Kosta & Roy, 2023; Negi et al., 2024), as detailed in Section 4.1, significantly lowers AEE on the DSEC-flow dataset (Table 1). Across both fully spiking and hybrid architectures, TSkips consistently yielded an average reduction in AEE of 14% and 9.5%, respectively. Notably, these improvements are achieved with models 3×3$\times$3 × smaller than baselines with comparable AEE. Although SOTA methods such as E-RAFT (Gehrig et al., 2021b) and E-FlowFormer (Li et al., 2023) achieve 6% and 10% lower AEE , respectively, these reductions come at the cost of increased complexity. E-RAFT, for example, while being 3×3$\times$3 × smaller than our best model, relies on complex processing of event data before its gated recurrent unit (GRU) update. These processing operations involve converting each event to a voxel grid, extracting features using CNNs and building a correlation volume. These dense kernel operations are computationally expensive during both training and inference, as they must be performed for every event sample. E-FlowFormer is a transformer-based architecture which relies on dense MAC operations in it’s self and cross attention blocks. These event processing steps lead to higher inference energy compared to TSkips architectures, which take advantage of the inherent sparsity and event-driven nature of SNNs, performing computationally cheaper sparse AC operations. Additionally, E-FlowFormer is trained on a large custom dataset and DSEC-flow, which introduces a significant increase in computation power and memory during training. Fig. 7 provides a qualitative analysis of how the 3×3$\times$3 × larger baselines perform compared to the smaller TSkips variants which highlights the efficiency of TSkips in allowing accurate temporal processing without increasing model complexity. Hyper parameters and a detailed analysis of convergence and inference energy for TSkips on DSEC-flow models are in Appendix B.
Table 1: AEE on DSEC-flow (lower is better). Best performing SNN and hybrid baselines and forward/backward TSkips models are highlighted in bold.
@@
EV-FlowNet (Gehrig et al., 2021b) 13.04 2.32 -
LSTM-FlowNet (Ponghiran et al., 2023) N/A [^1] N/A - not reported #Params 1.28 -
E-RAFT (Gehrig et al., 2021b) 5.3 0.79 -
E-FlowFormer (Li et al., 2023) N/A 0.76 -
@@
Base + Forward TSkips 1.12 0.86
Base + Backward TSkips 1.13 0.84
@@
Mini + Forward TSkips 1.44 1.02
Mini + Backward TSkips 1.46 1.01
@@
Micro + Forward TSkips 1.57 1.12
Micro + Backward TSkips 1.56 1.09
@@
Nano + Forward TSkips 1.86 1.37
Nano + Backward TSkips 1.77 1.35
@@
Pico + Forward TSkips 2.19 1.78
Pico + Backward TSkips 2.28 1.81
4.4 DVS128-Gesture, SHD and SSC Results
On the DVS Gesture dataset, incorporating TSkips resulted in a significant 8.37% improvement in accuracy over the baseline model, a convolutional SNN found by Kim et al. (2022). Importantly, our approach achieved this with significantly smaller models compared to SOTA, including a spiking transformer (Qin & Liu, 2024), demonstrating the efficiency of our proposed method. Furthermore, our method outperformed several other approaches, including spiking RNNs (Xing et al., 2020), methods incorporating learnable delays (Shrestha & Orchard, 2018), and methods using specialized LIF neurons or spatio-temporal feature extraction (Jiang & Zhang, 2024; Samadzadeh et al., 2023). This highlights the effectiveness of our approach in capturing and leveraging temporal dependencies for improved gesture recognition.
Table 2: Classification accuracy on DVS-Gesture. Comparisons between forward/backward TSkips and SOTA models. Best models in bold.
@@ 4.4 DVS128-Gesture, SHD and SSC Results
SLAYER (Shrestha & Orchard, 2018) SNN (8 layer) N/A 93.6
KLIF (Jiang & Zhang, 2024) VGG-11 N/A 93.75
DECOLE (Kaiser et al., 2020) Custom N/A 95.2
SNN-Skip (Benmeziane et al., 2023) ResNet18 N/A 95.43
Streaming rollout SNN Kugele et al. (2021) DenseNet 0.8 95.56
SCRNN (Xing et al., 2020) Custom N/A 96.59
STS-ResNet (Samadzadeh et al., 2023) ResNet18 N/A 96.7
Mamba-Spike (Qin & Liu, 2024) Mambda 6.1 97.8
@@
vRNN [^2] Backward skips are added to all baseline layers with $\Delta t=1$. 0.30 89.12
Forward TSkips 0.42 94.82
Backward TSkips 0.38 95.97
Forward + Backward TSkips 0.55 97.52
Table 3: Classification accuracy on SHD and SSC datasets. Comparisons between forward, backward and a combination of both TSkips against SOTA models and two baseline models: a shallow 4-layer MLP network (Baseline - 1) and a deep 8-layer MLP (Baseline - 2). Best models with a single and two TSkips in bold.
@@
LSTM-LIF (Zhang et al., 2023) 0.14 0.11 88.91 63.46
SRNN (Yin et al., 2021) N/A N/A 90.4 74.2
DL256-SNN-DLoss (Sun et al., 2023a) 0.14 - 92.56 -
SpikGRU (Dampfhoffer et al., 2022) - 0.28 - 77.00
radLIF (Bittar & Garner, 2022) 3.9 3.9 94.62 77.40
Spikingformer (Zhou et al., 2023) [^3] Zhou et al. (2023) was evaluated on SHD and SSC with minimal updates to the architecture and tuned learning rate. 1.98 1.99 82.68 72.43
DCLS-Delays (Sun et al., 2023b) 0.21 2.5 95.07 80.69
@@
vRNN22footnotemark: 2 - 1 0.16 0.12 68.50 71.20
Forward TSkips - 1 0.24 0.42 92.32 76.50
Backward TSkips - 1 0.19 0.24 93.64 79.87
Forward + Backward TSkips - 1 0.20 0.54 93.01 78.64
vRNN22footnotemark: 2 - 2 1.04 1.04 71.35 72.24
Forward TSkips - 2 1.12 1.08 94.15 78.98
Backward TSkips - 2 1.16 1.14 93.86 79.65
Forward + Backward TSkips - 2 1.28 1.42 94.73 80.23
@@
While TSkips achieves accuracy within 0.34% of SOTA model DCLS-Delays (Hammouamri et al., 2023) on SHD and 0.46% on SSC, it offers several distinct advantages. Specifically, the architectural constraint of DCLS-Delays to feedforward networks highlights a key area where TSkips provides a complementary alternative. Furthermore, TSkips as an architectural addition coupled with NAS allows for modifications to various network architectures, as shown in Section 4.3 and also offers efficiency due to very few new trainable parameters coupled with strong performance. In particular TSkips achieves accuracy comparable to Hammouamri et al. (2023) on the SSC dataset with a 44% smaller model.